GROWTH RATES Sawyer Corporation's 2015 sales were $5 million. Its 2010 sales were $2.5 million. At what rate have sales been growing? Round your answer to two decimal places. % Suppose someone made this statement: "Sales doubled in 5 years. This represents a growth of 100% in 5 years, so, dividing 100% by 5, we find the growth rate to be 20% per year." Is the statement correct?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

5=2.5(1+r/100)^5

(5/2.5)^(1/5)=(1+r/100)

(1+r/100)=1.1487

Hence r=(1.1487-1)*100

=14.87%(Approx)

Hence the statement is incorrect due to compounding effect ie 'interest earned on interest factor'.